Integrative Risk and valuation   

Giant​ Enterprises' stock has a required return of 16.7 ​%. The​ company, which plans to pay a dividend of ​$1.69 per share in the coming​ year, anticipates that its future dividends will increase at an annual rate consistent with that experienced over 2013 ​-2019 ​period, when the following dividends were​ paid

2019   $1.61
2018   $1.53
2017   $1.46
2016   $1.39
2015   $1.32
2014   $1.26
2013   $1.20

a.  If the​risk-freerate is 6%,what is the risk premium on​Giant'sstock?

b.  Using the​ constant-growth model, estimate the value of​Giant's stock.  

​(​Hint:Round the computed dividend growth rate to the nearest whole​percent.)

c.  Explain what ​effect,if​ any,a decrease in the risk premium would have on the value of​ Giant's stock.
Answer #1

a

Risk premium = required rate-risk free rate = 16.7-6 = 10.7%

b

Annual average growth rate
=((last value/First value)^(1/Time between 1st and last value)-1)*100
=((1.61/1.2)^(1/6)-1)*100
Annual Growth rate% = 5.0
As per DDM
Price = Dividend in 1 year/(cost of equity - growth rate)
Price = 1.69/ (0.167 - 0.05)
Price = 14.44

c

If risk premium decreases , price will increase as price is inversely proportional to required rate and required rate is directly proportional to risk premium